# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v2

@@ 8. Proof of Proposition 2
Proof of $I_{\nu_{1}\otimes\nu_{2}}(a)\geq J_{1}(a)$: The isoperimetric profile in our setup can be equivalently defined as the infimum of the perimeters of the sets with a given volume and smooth boundaries [6, pp. 32-33]. So, it suffices to consider an arbitrary set $A$ of measure $a$ with smooth boundary. Let $\phi_{n}:[0,\infty)\to[0,1]$ be a smooth function such that $\phi_{n}(0)=0$ and $\phi_{n}(t)=1$ for all $t\geq 1/n$. Let $f_{n}(\mathbf{x})=\phi_{n}(d(\mathbf{x},A))$, where $d$ denotes the geodesic distance on $\mathcal{M}_{1}\times\mathcal{M}_{2}$. Then, $f_{n}$ is smooth. Moreover, by the coarea formula,
$\lim_{n\to\infty}\mathbb{E}|\nabla f_{n}|=(\nu_{1}\otimes\nu_{2})^{+}(A).$
Hence, it suffices to prove that
$\lim_{n\to\infty}\mathbb{E}|\nabla f_{n}|\geq J_{1}(a).$
@@
$\mathbb{E}|\nabla f_{n}|$ $=\int\int\sqrt{|\nabla_{x}f_{n}|^{2}+|\nabla_{y}f_{n}|^{2}}{\text{\rm d}}\nu_{2}{\text{\rm d}}\nu_{1}$
@@
(8.2) $=\int\sqrt{|\nabla_{x}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)|^{2}+\left(\int_{0}^{1}\nu_{2}^{+}\{y:f_{n}(x,y)>t\}{\text{\rm d}}t\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
$\geq\int\sqrt{|\nabla_{x}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)|^{2}+\left(\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
$=\frac{a_{n}}{a}\int\sqrt{|\nabla_{x}\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)|^{2}+\left(\frac{a}{a_{n}}\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
(8.3) $\geq\frac{a_{n}}{a}\int\sqrt{|\nabla_{x}\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)|^{2}+I_{\nu_{2}}(\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y))^{2}}{\text{\rm d}}\nu_{1}(x)$
$\qquad-\frac{a_{n}}{a}\int|I_{\nu_{2}}(\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y))-\frac{a}{a_{n}}\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t|{\text{\rm d}}\nu_{1}(x)$
(8.4) $\geq\frac{a_{n}}{a}J_{1}(a)-\int|\frac{a_{n}}{a}I_{\nu_{2}}(\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y))-\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t|{\text{\rm d}}\nu_{1}(x),$
@@
By the dominated convergence theorem and the continuity of probability measures,
$\lim_{n\to\infty}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)=\nu_{2}(A_{x}),$
and
@@
$=\int_{0}^{1}I_{\nu_{2}}(\lim_{n\to\infty}\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t$
$=\int_{0}^{1}I_{\nu_{2}}(\nu_{2}(A_{x})){\text{\rm d}}t$
$=I_{\nu_{2}}(\nu_{2}(A_{x})),$
@@
We next prove $I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)\leq J_{1}(a)$. Let $\sigma:\mathcal{M}_{1}\to[0,1]$ be a feasible solution to $J_{1}(a)$, i.e., it is smooth and $\int_{\mathcal{M}_{1}}\sigma(x){\text{\rm d}}\nu_{1}(x)=a$. Let
$g(x)=F_{2}^{-1}\circ\sigma(x),$
and
@@
$\int_{\mathcal{M}_{1}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+|\nabla\sigma(x)|^{2}}{\text{\rm d}}\nu_{1}(x)$
$=\int_{\mathcal{M}_{1}}\sqrt{f_{2}(g(x))^{2}+f_{2}(g(x))^{2}|\nabla g(x)|^{2}}{\text{\rm d}}\nu_{1}(x)$
$=\int_{\mathcal{M}_{1}}\sqrt{1+|\nabla g(x)|^{2}}f_{2}(g(x)){\text{\rm d}}\nu_{1}(x)$
$=\left(\nu_{1}\otimes\tilde{\nu}_{2}\right)^{+}(A)\geq I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a),$
@@
Let
$\varphi(t)=\sigma\circ F_{1}^{-1}(t),$
and
$A:=\{(t,y)\in[0,1]\times\mathbb{R}:y\leq\varphi(t)=\sigma\circ F_{1}^{-1}(t)\}.$
By definition,
@@
$\int_{\mathbb{R}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+(\sigma^{\prime}(x))^{2}}{\text{\rm d}}F_{1}(x)$
$=\int_{\mathbb{R}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+(\varphi^{\prime}(F_{1}(x))f_{1}(x))^{2}}{\text{\rm d}}F_{1}(x)$
@@ 9. Proof of Proposition 3
Assume that Conjecture 2 holds. Then, for any probability measures $\nu_{1},...,\nu_{k}$ satisfying (1.5), it holds that
@@
Let $A$ be an isoperimetric minimizer of probability $e^{-n\alpha}$ in the classic sense. The proof of Theorem 10 up to (3.29) still works for this case. Recall $\tilde{A}$ defined in (3.29), whose probability satisfies
$e^{-n\alpha^{\prime}}\leq\nu_{n}(\tilde{A})\leq e^{-n\alpha}$
with $\alpha^{\prime}:=\alpha+\delta^{\prime}$. Here, $\delta^{\prime}=\delta+H(\epsilon^{\prime})$ can be chosen arbitrarily small. Note that here $\nu_{n}:=\nu^{\otimes n}$.
@@
$\to\inf_{0\leq\alpha_{0}\leq\alpha_{\tau}}\breve{\Theta}(\alpha_{0})+2K_{1}(\alpha_{\tau}-\alpha_{0})$
$=\breve{\Theta}(\alpha_{\tau}),$
where $\alpha_{\tau}:=\liminf_{n\to\infty}\alpha_{n,\tau}$, $K_{1}:=K_{IS,1}^{+}=\lim_{a\to 0}\frac{I_{1}(a)}{a\sqrt{2\ln\frac{1}{a}}}$, and the last line follows since from Statement 3 of Theorem 2, it is known that $K_{1}\geq K_{IS}^{+}\geq K_{LS}^{+}\geq\frac{\breve{\Theta}^{\prime}(\alpha_{\tau})}{2}$.
@@
(9.3) $\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}\geq\sqrt{\breve{\Theta}(\alpha_{\tau})}-\frac{\delta^{\prime}}{\sqrt{\tau}}.$
@@
That is,
$\alpha_{\tau}\geq\alpha-(\sqrt{\tau\breve{\Theta}(\alpha)}+\delta^{\prime}).$
Substituting this into (9.3) yields that
(9.4) $\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}\geq\sqrt{\breve{\Theta}\left(\alpha-(\sqrt{\tau\breve{\Theta}(\alpha)}+\delta^{\prime})\right)}-\frac{\delta^{\prime}}{\sqrt{\tau}}.$
Letting $\delta^{\prime}\to 0$ first and $\tau\to 0$ then,
$\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}\geq\sqrt{\breve{\Theta}(\alpha)}.$

### v3

@@ Appendix B-1. Proof of Theorem 18
Proof of $I_{\nu_{1}\otimes\nu_{2}}(a)\geq J_{1}(a)$: The isoperimetric profile in our setup can be equivalently defined as the infimum of the perimeters of the sets with a given volume and smooth boundaries [10, pp. 32-33]. So, it suffices to consider an arbitrary set $A$ of measure $a$ with smooth boundary. Let $\phi_{n}:[0,\infty)\to[0,1]$ be a smooth function such that $\phi_{n}(0)=0$ and $\phi_{n}(t)=1$ for all $t\geq 1/n$. Let $f_{n}(\mathbf{x})=\phi_{n}(d(\mathbf{x},A))$, where $d$ denotes the geodesic distance on $\mathcal{M}_{1}\times\mathcal{M}_{2}$. Then, $f_{n}$ is smooth. Moreover, by the coarea formula, $\lim_{n\to\infty}\mathbb{E}|\nabla f_{n}|=(\nu_{1}\otimes\nu_{2})^{+}(A).$ Hence, it suffices to prove that $\lim_{n\to\infty}\mathbb{E}|\nabla f_{n}|\geq J_{1}(a).$
@@
$\quad\mathbb{E}|\nabla f_{n}|=\int\int\sqrt{|\nabla_{x}f_{n}|^{2}+|\nabla_{y}f_{n}|^{2}}{\text{\rm d}}\nu_{2}{\text{\rm d}}\nu_{1}$
@@
(6.29) $=\int\sqrt{\Big|\nabla_{x}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)\Big|^{2}+\left(\int_{0}^{1}\nu_{2}^{+}\{y:f_{n}(x,y)>t\}{\text{\rm d}}t\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
$\geq\int\sqrt{\Big|\nabla_{x}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)\Big|^{2}+\left(\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
$=\frac{a_{n}}{a}\int\sqrt{\Big|\nabla_{x}\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)\Big|^{2}+\left(\frac{a}{a_{n}}\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\right)^{2}}{\text{\rm d}}\nu_{1}(x)$
(6.30) $\geq\frac{a_{n}}{a}\int\sqrt{\Big|\nabla_{x}\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)\Big|^{2}+I_{\nu_{2}}(\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y))^{2}}{\text{\rm d}}\nu_{1}(x)$
$\qquad-\frac{a_{n}}{a}\int\Big|I_{\nu_{2}}(\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y))-\frac{a}{a_{n}}\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\Big|{\text{\rm d}}\nu_{1}(x)$
(6.31) $\geq\frac{a_{n}}{a}J_{1}(a)-\int\Big|\frac{a_{n}}{a}I_{\nu_{2}}(\frac{a}{a_{n}}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y))-\int_{0}^{1}I_{\nu_{2}}(\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t\Big|{\text{\rm d}}\nu_{1}(x),$
@@
By the dominated convergence theorem and the continuity of probability measures, $\lim_{n\to\infty}\int f_{n}(x,y){\text{\rm d}}\nu_{2}(y)=\nu_{2}(A_{x}),$ and
@@
$\qquad=\int_{0}^{1}I_{\nu_{2}}(\lim_{n\to\infty}\nu_{2}\{y:f_{n}(x,y)>t\}){\text{\rm d}}t=\int_{0}^{1}I_{\nu_{2}}(\nu_{2}(A_{x})){\text{\rm d}}t=I_{\nu_{2}}(\nu_{2}(A_{x})),$
@@
We next prove $I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)\leq J_{1}(a)$. Let $\sigma:\mathcal{M}_{1}\to[0,1]$ be a feasible solution to $J_{1}(a)$, i.e., it is smooth and $\int_{\mathcal{M}_{1}}\sigma(x){\text{\rm d}}\nu_{1}(x)=a$. Let $g(x):=F_{2}^{-1}\circ\sigma(x),$ and
@@
$\int_{\mathcal{M}_{1}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+|\nabla\sigma(x)|^{2}}{\text{\rm d}}\nu_{1}(x)=\int_{\mathcal{M}_{1}}\sqrt{f_{2}(g(x))^{2}+f_{2}(g(x))^{2}|\nabla g(x)|^{2}}{\text{\rm d}}\nu_{1}(x)$
$\qquad=\int_{\mathcal{M}_{1}}\sqrt{1+|\nabla g(x)|^{2}}f_{2}(g(x)){\text{\rm d}}\nu_{1}(x)=\left(\nu_{1}\otimes\tilde{\nu}_{2}\right)^{+}(A)\geq I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a),$
@@
Let $\varphi:=\sigma\circ F_{1}^{-1}$. By definition,
@@
$\int_{\mathbb{R}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+(\sigma^{\prime}(x))^{2}}{\text{\rm d}}F_{1}(x)$ $=\int_{\mathbb{R}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+(\varphi^{\prime}(F_{1}(x))f_{1}(x))^{2}}{\text{\rm d}}F_{1}(x)$
@@
Assume that Conjecture 2 holds. Then, for any probability measures $\nu_{1},...,\nu_{k}$ satisfying Assumption 1, it holds that
@@
Let $A$ be an isoperimetric minimizer of probability $e^{-n\alpha}$ in the standard sense. The proof of Theorem 10 up to (3.29) still works for this case. Recall $\tilde{A}$ defined in (3.29), whose probability satisfies $e^{-n\alpha^{\prime}}\leq\nu_{n}(\tilde{A})\leq e^{-n\alpha}$ with $\alpha^{\prime}:=\alpha+\delta^{\prime}$. Here, $\delta^{\prime}=\delta+H(\epsilon^{\prime})$ can be chosen arbitrarily small. Note that here $\nu_{n}:=\nu^{\otimes n}$.
@@
$\to\inf_{0\leq\alpha_{0}\leq\alpha_{\tau}}\breve{\Theta}(\alpha_{0})+2K_{1}(\alpha_{\tau}-\alpha_{0})\,=\breve{\Theta}(\alpha_{\tau}),$
where $\alpha_{\tau}:=\liminf_{n\to\infty}\alpha_{n,\tau}$, $K_{1}:=K_{IS,1}^{+}=\lim_{a\to 0}\frac{I_{1}(a)}{a\sqrt{2\ln\frac{1}{a}}}$, and the equality follows since from Statement 3 of Theorem 2, it is known that $K_{1}\geq K_{IS}^{+}\geq K_{LS}^{+}\geq\frac{\breve{\Theta}^{\prime}(\alpha_{\tau})}{2}$.
@@
(6.34) $\underline{\Lambda}(\alpha)\geq\sqrt{\breve{\Theta}(\alpha_{\tau})}-\frac{\delta^{\prime}}{\sqrt{\tau}}.$
@@
That is, $\alpha_{\tau}\geq\alpha-(\sqrt{\tau\breve{\Theta}(\alpha)}+\delta^{\prime}).$ Substituting this into (6.34) yields that $\underline{\Lambda}(\alpha)\geq\sqrt{\breve{\Theta}\left(\alpha-(\sqrt{\tau\breve{\Theta}(\alpha)}+\delta^{\prime})\right)}-\frac{\delta^{\prime}}{\sqrt{\tau}}.$ Letting $\delta^{\prime}\to 0$ first and $\tau\to 0$ then, $\underline{\Lambda}(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}.$